A netforce of 16 N causes a mass to accelerate at a rate of 5 m/s2. Determine the mass.
1
Expert's answer
2020-11-23T10:26:21-0500

We can find the mass from the Newton's Second Law of Motion:


"F_{net}=ma,""m=\dfrac{F_{net}}{a}=\dfrac{16\ N}{5\ \dfrac{m}{s^2}}=3.2\ kg."

Answer:

"m=3.2\ kg."